# Resurgence in Lorentzian quantum cosmology: No-boundary saddles and resummation of quantum gravity corrections around tunneling saddle points

### v2

@@ I Introduction
To avoid these problems, the authors of Ref. [6] suggested path integrals along the steepest descent paths in complex metrics. This method does not rely on starting with Euclidean or Lorentzian metrics. Instead, by treating it as complex, integrals are done along contours where the real part of the action decreases rapidly. However, these contours are not unique, causing ambiguity between the Hartle-Hawking no-boundary proposal [2] and Vilenkin’s tunneling proposal [7] (related early proposals were given in Refs. [8, 9, 10, 11]). These proposals lead to the wave function of the Universe with the opposite exponential dependence, $G\sim\exp(\pm 24\pi^{2}/\hbar\Lambda)$ , where the $+$ and -- sign correspond to the no-boundary and tunneling proposal, respectively. Depending on the sign, physical consequences are drastically different [12, 13].
@@ IV Neumann boundary condition and airy function
Although the Lefschetz thimble structure and resurgence analysis of the Airy functions are well known, by examining these analytical properties of the Airy functions, we can re-evaluate the wave function in Lorentzian quantum cosmology. Now, by transforming $\bar{N}\to\bar{N}\sqrt{\hbar}$, we have the following expression,
$\Psi(\hbar)=\int_{-\infty}^{\infty}\mathrm{d}\bar{N}\exp\left(i\bar{\alpha}\hbar^{2}\bar{N}^{3}+i\bar{\beta}\bar{N}\right)=\frac{2\pi\text{Ai}\left(\frac{\bar{\beta}}{\sqrt[3]{3\bar{\alpha}\hbar^{2}}}\right)}{\sqrt[3]{3\bar{\alpha}\hbar^{2}}}\,,$ (33)
@@
$\begin{split}c_{n}=&\frac{(3\bar{\alpha}\bar{\beta}\hbar^{2})^{-\frac{1}{4}}}{2}e^{-\frac{2}{3\hbar}\sqrt{\frac{\bar{\beta}^{3}}{3\bar{\alpha}}}}\left((-1)^{3n}+1\right)\\ &\times\frac{\Gamma\left(\frac{3n}{2}+\frac{1}{2}\right)}{\Gamma(n+1)}\left(i\sqrt[4]{\frac{\bar{\alpha}}{27\bar{\beta}^{3}}}\right)^{n}.\end{split}$ (36)
@@
${\cal B}\Psi_{+}(t)=\sqrt[4]{\frac{\pi^{2}}{3\bar{\alpha}\bar{\beta}\hbar^{2}}}e^{-\frac{2}{3\hbar}\sqrt{\frac{\bar{\beta}^{3}}{3\bar{\alpha}}}}\,_{2}F_{1}\left(\frac{1}{6},\frac{5}{6};1;-\frac{3}{4}t\sqrt{\frac{3\bar{\alpha}}{\bar{\beta}^{3}}}\right)\,.$ (37)
@@
$\begin{split}\mathcal{S}_{\theta}\Psi_{+}(\hbar)&=\frac{1}{\hbar}\int_{0}^{\infty\cdot e^{i\theta}}\mathrm{d}t\,e^{-\frac{t}{\hbar}}{\cal B}\Psi_{+}(t)\\ &=\frac{2\pi\text{Ai}\left(\frac{\bar{\beta}}{\sqrt[3]{3\bar{\alpha}\hbar^{2}}}\right)}{\sqrt[3]{3\bar{\alpha}\hbar^{2}}}\,,\end{split}$ (38)

### v3

@@ I Introduction
To avoid these problems, the authors of Ref. [6] suggested path integrals along the steepest descent paths in complex metrics. This method does not rely on starting with Euclidean or Lorentzian metrics. Instead, by treating it as complex, integrals are done along contours where the real part of the action decreases rapidly. However, these contours are not unique, causing ambiguity between the Hartle-Hawking no-boundary proposal [2] and Vilenkin’s tunneling proposal [7] (related early proposals were given in Refs. [8, 9, 10, 11]). These proposals lead to the wave function of the Universe with the opposite exponential dependence, $G\sim\exp(\pm 12\pi^{2}/\hbar\Lambda)$ , where the $+$ and -- sign correspond to the no-boundary and tunneling proposal, respectively. Depending on the sign, physical consequences are drastically different [12, 13].
@@ IV Neumann boundary condition and airy function
Although the Lefschetz thimble structure and resurgence analysis of the Airy functions are well known, by examining these analytical properties of the Airy functions, we can re-evaluate the wave function in Lorentzian quantum cosmology. Now, by transforming $\bar{N}\to\bar{N}\hbar$, we have the following expression,
$\Psi(\hbar)=\hbar\int_{-\infty}^{\infty}\mathrm{d}\bar{N}\exp\left(i\bar{\alpha}\hbar^{2}\bar{N}^{3}+i\bar{\beta}\bar{N}\right)=\frac{2\pi\hbar^{\frac{1}{3}}\text{Ai}\left(\frac{\bar{\beta}}{\sqrt[3]{3\bar{\alpha}\hbar^{2}}}\right)}{\sqrt[3]{3\bar{\alpha}}}\,,$ (33)
@@
$\begin{split}c_{n}=&\frac{\sqrt{\hbar}}{2(3\bar{\alpha}\bar{\beta})^{\frac{1}{4}}}e^{-\frac{2}{3\hbar}\sqrt{\frac{\bar{\beta}^{3}}{3\bar{\alpha}}}}\left((-1)^{n}+1\right)\\ &\times\frac{\Gamma\left(\frac{3n}{2}+\frac{1}{2}\right)}{\Gamma(n+1)}\left(i\sqrt[4]{\frac{\bar{\alpha}}{27\bar{\beta}^{3}}}\right)^{n}.\end{split}$ (36)
@@
${\cal B}\Psi_{+}(t)=\sqrt[4]{\frac{\pi^{2}\hbar^{2}}{3\bar{\alpha}\bar{\beta}}}e^{-\frac{2}{3\hbar}\sqrt{\frac{\bar{\beta}^{3}}{3\bar{\alpha}}}}\,_{2}F_{1}\left(\frac{1}{6},\frac{5}{6};1;-\frac{3}{4}t\sqrt{\frac{3\bar{\alpha}}{\bar{\beta}^{3}}}\right)\,.$ (37)
@@
$\begin{split}\mathcal{S}_{\theta}\Psi_{+}(\hbar)&=\frac{1}{\hbar}\int_{0}^{\infty\cdot e^{i\theta}}\mathrm{d}t\,e^{-\frac{t}{\hbar}}{\cal B}\Psi_{+}(t)\\ &=\frac{2\pi\hbar^{\frac{1}{3}}\text{Ai}\left(\frac{\bar{\beta}}{\sqrt[3]{3\bar{\alpha}\hbar^{2}}}\right)}{\sqrt[3]{3\bar{\alpha}}}\,,\end{split}$ (38)